# FedRTS: Federated Robust Pruning via Combinatorial Thompson Sampling

### v1

@@ F.1 Compression Strategy
  • For density ($d\in[0.9,1]$), a dense scheme is utilized, leading to $s=n\textperiodcentered b$.
@@
F.2 Communication Cost
In terms of communication costs, FedRTS shares a communication cost strategy similar to that of FedTiny. In contrast to other baseline methods like FedDST, where clients need to upload TopK gradients to the server every $\Delta R$ rounds to facilitate parameter expansion, the quantity of TopK gradients, denoted as $\xi_{t}$, aligns with the number of marked parameters $\theta$ low. Here, $\xi_{t}=\zeta_{t}(1-s_{m})n_{\theta}$, with $\zeta_{t}=0.2(1+cos{\frac{t\pi}{R_{stop}E}})$representing the adjustment rate for the t-th iteration. This results in minimal upload overhead. Moreover, there is no communication overhead for the model mask m during download since sparse storage formats like bitmap and coordinate offset contain the same element position information. Auxiliary data such as the learning rate schedule is excluded.
We define the storage for dense and sparse parameters as $O_{d}$ and $O_{s}$ respectively. Notably, in the inner loop, all federated pruning methods except FedAVG share the same communication cost, which amounts to $2O_{s}$. However, in the outer loop, different federated learning frameworks exhibit varying communication costs. The communication costs of different federated learning frameworks in the outer loop are detailed below:
@@ F.2 Communication Cost
  • FedSGC: FedSGC requires the server to send the gradient congruity to the server for each round, which has the same size with sparse weight. Therefore, the data exchange in the outer loop is $3O_{s}$.
@@ F.3 Computational FLOPs
[…]

### v3

@@ G.4.1 Compression Strategy
  • For density ($d\in[0.9,1]$), a dense scheme is utilized, leading to $s=n·b$.
@@
G.4.2 Communication Cost
In terms of communication costs, FedRTS shares a communication cost strategy similar to that of FedTiny. In contrast to other baseline methods like FedDST, where clients need to upload TopK gradients to the server every $\Delta R$ rounds to facilitate parameter expansion, the quantity of TopK gradients, denoted as $\xi_{t}$, aligns with the number of marked parameters $\theta$ low. Here, $\xi_{t}=\zeta_{t}(1-s_{m})n_{\theta}$, with $\zeta_{t}=0.2(1+cos{\frac{t\pi}{R_{stop}E}})$representing the adjustment rate for the t-th iteration. This results in minimal upload overhead. Moreover, there is no communication overhead for the model mask m during download since sparse storage formats like bitmap and coordinate offset contain the same element position information. Auxiliary data such as the learning rate schedule are excluded.
We define the storage for dense and sparse parameters as $O_{d}$ and $O_{s}$, respectively. Notably, in the inner loop, all federated pruning methods except FedAVG share the same communication cost, which amounts to $2O_{s}$. However, in the outer loop, different federated learning frameworks exhibit varying communication costs. The communication costs of different federated learning frameworks in the outer loop are detailed below:
@@ G.4.2 Communication Cost
  • FedSGC: FedSGC requires the server to send the gradient congruity to the server for each round, which has the same size as sparse weight. Therefore, the data exchange in the outer loop is $3O_{s}$.
@@ G.4.3 Computational FLOPs
In a detailed breakdown, assuming inference FLOPs for dense and static sparse models are denoted as $F_{d}$ and $F_{s}$ respectively, and the local iteration count is $E$, the maximum training FLOPs for each framework are as follows: